OpenStudy (anonymous):

list the first 20 counting numbers in base 6 ???????????????????????????????????????????????????????????????????????????????????????????

OpenStudy (anonymous):

1 2 3 4 5 10 11 12 13 14 20 21 22 23 24 25 30 31 32 33

OpenStudy (anonymous):

How did you get that answer? what did u do?

OpenStudy (anonymous):

There are only 6 possible digits in a base 6 system, 0, 1, 2, 3, 4, 5. Once you get to the last one (5), you must use the next place to the left, so the next 6 numbers are 10, 11, 12, 13, 14, 15. Then the place value to the left must be increased again to 2, making the next six, 20, 21, 22, 23, 24, 25. Hope that helps

OpenStudy (anonymous):

Instead of the place values being the ones, tens, hundreds, etc. that you would find in the base 10 or decimal system, the place values are the ones, sixes, thirty-sixes, etc.

OpenStudy (anonymous):

in base 6 there are only 6 1 digit numbers :0 1 2 3 4 5: 10 = 6 + 1 = 7 in base 10 20 = 2(6) = 12 in base 10 33 = 3(6) + 3 = 21 in base 10

OpenStudy (anonymous):

please explain the "base" concept. perhaps then i will understand

OpenStudy (anonymous):

You could think of the base system that you are probably most familiar with, base ten. You count, 0, 1, 2, 3, 4, 5, 6, 7, 8, 9 (total of 10 digits in the base 10 system). When you get beyond 9, you need another place to accommodate the greater numbers, so you include the 10's place. One ten is equal to 10 ones. 10 actually means 1 ten and 0 ones. In the base 6 system, you can't have a single digit greater than 5. In order to make greater numbers, you must include another place, so the "ones" reset back to zero, and in this case, you have 1 six and 0 ones or 10 (which would actually mean 6 in the base ten system.

OpenStudy (anonymous):

In hexadecimal (base 16) which you might have also come across, there are also 6 letters included, so you count 0, 1, 2, 3, 4, 5, 6, 7, 8, 9, A, B, C, D, E, F, 10, 11, etc.